OpenStudy (anonymous):

x/5 - y/6 = 1 can I graph this?

OpenStudy (anonymous):

Of course YES.....

OpenStudy (anonymous):

not sure how..do I place the number 5 on the x axis

OpenStudy (anonymous):

Its in the double-intercept from

OpenStudy (anonymous):

actually 6 on the x axis and 5 on the y axis

OpenStudy (anonymous):

it might be easier to graph \(6x-5y=30\)

hero (hero):

Hint: to graph, add y/6 to both sides, subtract one from both sides: x/5 - 1 = y/6 Now multiply both sides by 6 to get 6x/5 - 6 = y

hero (hero):

Now you might be able to graph that

OpenStudy (anonymous):

There,